OpenStudy (anonymous):

Suppose f(π/3) = 4 and f '(π/3) = −7, and let g(x) = f(x) sin x and h(x) = (cos x)/f(x). Find the following. a. g'(π/3) b. h'(π/3)

zepdrix (zepdrix):

\[\Large\bf\sf h(x)\quad=\quad \frac{\cos x}{f(x)}\] \[\Large\bf\sf h'(x)\quad=\quad \frac{\color{royalblue}{\left[\cos x\right]'}f(x)-\cos x\color{royalblue}{\left[f(x)\right]'}}{f(x)^2}\]We need to apply quotient rule for this problem. Understand how I set that up? We need to take the derivative of the blue parts.

OpenStudy (anonymous):

Can you elaborate?